the area of a rectangle is 6800. if the width of the feild is 80 meters, what is the perimeter of the feild?
Answer by Kitneaadmeethe(2)

Given,Area=6800
and width,w=80
let 'l' be the length of the rectangle.
Now we know that area of rectangle=length*width
therefore area=l*80
but area is already given
i.e. Area=l*80=6800(given)
l=6800/80
l=85
now,we have both length and width
now we know that perimeter is the total outline covering any particular diagram,in case of an rectangle we have a cover of two lengths and two widths i.e.
Perimeter=2l+2w
perimeter=2(l+w)
perimeter=2(85+80)
perimeter=2*165
perimeter=330 ans.